# The Arens–Michael envelope of a solvable Lie algebra is a homological epimorphism

### v1

@@ 2. Hopf $\mathop{\widehat{\otimes}}$-algebras, analytic smash products and Arens–Michael envelopes
Recall that an Arens–Michael algebra is a topological algebra whose topology can be determined by a family of submultiplicative seminorms. The Arens–Michael envelope of a topological $\mathbb{C}$-algebra $A$ is a pair $(\widehat{A},\iota_{A})$, where $\widehat{A}$ is an Arens–Michael algebra and $\iota_{A}$ is a continuous homomorphism $A\to\widehat{A}$ such that for every Arens–Michael algebra $B$ and every continuous homomorphism $\varphi\!:A\to B$ there is a unique continuous homomorphism $\widehat{\varphi}\!:\widehat{A}\to B$ making the diagram
@@
commutative. More concrete, we can take as $\iota_{A}$ the completion homomorphism with respect to the topology determined by all possible (continuous) submultiplicative seminorms on $A$.
In the case when $A$ is an arbitrary associative $\mathbb{C}$-algebra, we can endow it with the strongest locally convex topology or, equivalently, take arbitrary homomorphism $\varphi$ in the definition. In the Lie algebra case, we can replace $U(\mathfrak{g})$ by $\mathfrak{g}$ in the diagram (assuming that $\varphi$ is a Lie algebra homomorphism) and say that $\widehat{U}(\mathfrak{g})$ is the Arens–Michael envelope of $\mathfrak{g}$.
@@
Let $\mathfrak{g}$ be a finite-dimensional solvable complex Lie algebra and fix an iterated semidirect sum decomposition $\mathfrak{g}=((\cdots(\mathfrak{f}_{1}\rtimes\mathfrak{f}_{2})\rtimes\cdots)\rtimes\mathfrak{f}_{n}$, where $\mathfrak{f}_{1},\ldots,\mathfrak{f}_{n}$ are $1$-dimensional. Then there is a non-increasing sequence $i_{1},\ldots,i_{n}$ in $[0,\infty]$ such that
@@
We use reduction to algebras of analytic functionals. Let $G$ be a simply connected complex Lie group whose Lie algebra is isomorphic to $\mathfrak{g}$. Also, let $\mathscr{A}(G)$ denote the corresponding algebra of analytic functionals and $\widehat{\mathscr{A}}(G)$ the Arens–Michael algebra of $\mathscr{A}(G)$. Then the natural embedding $U(\mathfrak{g})\to\mathscr{A}(G)$ induces a topological isomorphism $\widehat{U}(\mathfrak{g})\to\widehat{\mathscr{A}}(G)$ [2, Proposition 2.1]. So we can apply results on $\widehat{\mathscr{A}}(G)$, Theorems 6.5 and 6.4 in [9], which give the existence of the decomposition in (2.8) and the compatibility with the decomposition of $U(\mathfrak{g})$, respectively. ∎
@@
We use the results in this section when proving Theorem 3.4.
@@
$h\cdot\varphi\alpha(r)=\sum\psi(h_{(1)})\,\varphi\alpha(r)\,\psi(S(h_{(2)}))=\chi\left(\sum j(h_{(1)})\,i_{R}(r)\,j(S(h_{(2)}))\right)=\chi i_{R}(h\cdot r).$
(The last equality follows from the fact that $i_{R}$ is an $H$$\mathbin{\widehat{\otimes}}$-module morphism with respect to the adjoint action.) Thus $h\cdot\varphi\alpha(r)=\varphi\alpha(h\cdot r)$. Hence $\varphi\alpha$ is an $H$$\mathbin{\widehat{\otimes}}$-module morphism.
Further, by the hypothesis, $\alpha$ is also an $H$$\mathbin{\widehat{\otimes}}$-module morphism and so is $\varphi$ since the range of $\alpha$ is dense. Hence, by the universal property of smash product written for $S\mathop{\widehat{\#}}H$, there is a unique $\mathop{\widehat{\otimes}}$-algebra homomorphism $\theta$ such that $\theta i_{S}=\varphi$ and $\theta(\alpha\mathop{\widehat{\#}}1)j=\psi$.
Since $\varphi\alpha$ is an $H$$\mathbin{\widehat{\otimes}}$-module algebra homomorphism, it follows from the universal property of smash product written for $R\mathop{\widehat{\#}}H$ that a homomorphism $\chi$ such that $\chi i_{R}=\varphi\alpha$ and $\chi j=\psi$ is unique. Therefore $\theta(\alpha\mathop{\widehat{\#}}1)=\chi$.
@@
Similarly, using density again, we have that $(1\mathop{\widehat{\#}}\beta)j_{H}$ is a $K$$\mathbin{\widehat{\otimes}}$-module morphism with respect to the adjoint action associated with $j_{K}$. Hence, by the universal property of smash product written for $S\mathop{\widehat{\#}}K$, there is a unique $\mathop{\widehat{\otimes}}$-algebra homomorphism $\theta$ such that $\theta j_{K}=\psi$ and $\theta(1\mathop{\widehat{\#}}\beta)i=\varphi$. The first equality implies that $\theta(1\mathop{\widehat{\#}}\beta)j_{H}=\psi\beta$.
Being a $K$$\mathbin{\widehat{\otimes}}$-module morphism, $\varphi$ is also an $H$$\mathbin{\widehat{\otimes}}$-module morphism. Then the universal property of smash product written for $S\mathop{\widehat{\#}}H$ implies that a homomorphism $\chi$ such that $\chi i=\varphi$ and $\chi j_{H}=\psi\beta$ is unique. Therefore $\theta(1\mathop{\widehat{\#}}\beta)=\chi$.
@@ 3. Unique extension property for derivations
When $A$ is a $\mathop{\widehat{\otimes}}$-algebra and $X$ is an $A$$\mathop{\widehat{\otimes}}$-bimodule, we denote the vector space of continuous derivations from $A$ to $X$ by $\operatorname{Der}(A,X)$. It is clear that a $\mathop{\widehat{\otimes}}$-algebra homomorphism $\varphi\!:A\to B$ induces the linear map
$\widetilde{\varphi}_{X}\!:\operatorname{Der}(B,X)\to\operatorname{Der}(A,X)\!:D\mapsto Df.$
@@
It is proved in [28] that the Arens–Michael envelope always satisfies to Property (UDE).
In what follows we denote the pushout of $\mathop{\widehat{\otimes}}$-algebra homomorphisms $A\to B_{1}$ and $A\to B_{2}$ using relative free algebra notation, i.e., as $B_{1}\mathbin{\ast_{A}}B_{2}$. Note that the category of $\mathop{\widehat{\otimes}}$-algebras is cocomplete and, in particular, pushout exists. But we do not need this fact because we use only Propositions 2.6 and 2.7, which in particular imply the existence of pushouts.
The following result is an analytic version of Proposition 5.2 in [14].
@@
In the category of unital $\mathop{\widehat{\otimes}}$-algebras, Property (UDE) is preserved by pushouts.
For the proof we need an auxiliary proposition. Note that every $(B_{1}\mathbin{\ast_{A}}B_{2})$$\mathop{\widehat{\otimes}}$-bimodule is both a $B_{1}$$\mathop{\widehat{\otimes}}$- and $B_{2}$$\mathop{\widehat{\otimes}}$-bimodule with the multiplications given by lifting along homomorphisms.
@@
Let $\varphi_{1}\!:A\to B_{1}$ and $\varphi_{2}\!:A\to B_{2}$ be homomorphisms of $\mathop{\widehat{\otimes}}$-algebras and $X$ be a $(B_{1}\mathbin{\ast_{A}}B_{2})$$\mathop{\widehat{\otimes}}$-bimodule. Suppose that $D_{1}\in\operatorname{Der}(B_{1},X)$, $D_{2}\in\operatorname{Der}(B_{2},X)$ and $D_{1}\varphi_{1}=D_{2}\varphi_{2}$. Then there exists a unique continuous derivation $D$ making the diagram
@@
$(c,x)(c^{\prime},y)=(cc^{\prime},cy+xc^{\prime})\qquad(c,c^{\prime}\in B_{1}\mathbin{\ast_{A}}B_{2},\;x,y\in X).$
@@
Let
[…]
be a pushout diagram such that $\varphi_{1}$ satisfies Property (UDE).
Take a $(B_{1}\ast_{A}B_{2})$$\mathop{\widehat{\otimes}}$-bimodule $X$ and $D_{2}\in\operatorname{Der}(B_{2},X)$. Since $D_{2}\varphi_{2}$ is a derivation of $A$ and $\varphi_{1}$ satisfies Property (UDE), there exists $D_{1}\in\operatorname{Der}(B_{1},X)$ such that $D_{1}\varphi_{1}=D_{2}\varphi_{2}$. By the universal property for derivations in Proposition 3.3, there is $D\in\operatorname{Der}(B_{1}\ast_{A}B_{2},X)$ such that $D_{1}=D\varkappa_{1}$ and $D_{2}=D\varkappa_{2}$. Hence, $\operatorname{Der}(B_{1}\ast_{A}B_{2},X)\to\operatorname{Der}(B_{2},X)$ is surjective.
Note that $\widetilde{\varphi}_{X}$ is injective for each $X$ if and only if $\varphi$ is an epimorphism; cf. the Fréchet algebra case in [10, Theorem 3.20]. Since epimorphisms are preserved by pushouts in an arbitrary category, $\operatorname{Der}(B_{1}\ast_{A}B_{2},X)\to\operatorname{Der}(B_{2},X)$ is injective for each $X$. Thus, $\varkappa_{2}$ satisfies Property (UDE). ∎
@@
It is easy to see that Property (UDE) is stable under composition. So it suffices to show that $\alpha\mathop{\widehat{\#}}1$ and $1\mathop{\widehat{\#}}\beta$ have this property. Since it is preserved by pushouts according to Proposition 3.2, the result follows from Propositions 2.6 and 2.7, which assert that $\alpha\mathop{\widehat{\#}}1$ and $1\mathop{\widehat{\#}}\beta$ are obtained by pushouts with $\alpha$ and $\beta$, respectively. ∎
@@ 4. Homological epimorphisms and relatively quasi-free algebras
Suppose that $R$ is a $\mathop{\widehat{\otimes}}$-algebra. Recall that an $R$$\mathop{\widehat{\otimes}}$-algebra is a pair $(A,\eta_{A})$, where $A$ is a $\mathop{\widehat{\otimes}}$-algebra and $\eta_{A}\!:R\to A$ is a $\mathop{\widehat{\otimes}}$-algebra homomorphism. Note that each $A$$\mathop{\widehat{\otimes}}$-module is automatically an $R$$\mathop{\widehat{\otimes}}$-module via the restriction-of-scalars functor along $\eta_{A}$. We denote by $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}$ the category of left $A$$\mathop{\widehat{\otimes}}$-modules endowed with the split exact structure. This means that the admissible sequences in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}$ are those that split by $R$$\mathop{\widehat{\otimes}}$-module morphisms; cf. [28, Appendix, Example 10.1 and 10.3]. In particular, when $R=\mathbb{C}$, we recover the standard definition of an admissible (or $\mathbb{C}$-split) sequence of $A$$\mathop{\widehat{\otimes}}$-modules used in [18]. When considering $\mathop{\widehat{\otimes}}$-bimodules over an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ (from the left) and an $S$$\mathop{\widehat{\otimes}}$-algebra $B$ (from the right), we denote the corresponding category by $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(B,S)$. In the case when $R=S=\mathbb{C}$, we write simply $A\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}B$. When we talk on a projective resolution in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(B,S)$, we mean a complex consisting of objects projective in the corresponding exact category and splitting in $S\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}R$.
Let $A$ be a $\mathop{\widehat{\otimes}}$-algebra and let $X$ and $Y$ be a right and left $A$$\mathop{\widehat{\otimes}}$-modules, respectively. Then the projective $A$-module tensor product $X\mathop{\widehat{\otimes}}_{A}Y$ is defined as the completion of the quotient space of $X\mathop{\widehat{\otimes}}Y$ by the closure of the linear hull of all elements of the form $x\cdot a\otimes y-x\otimes a\cdot y$ ($x\in X$, $y\in Y$, $a\in A$); see [18].
@@
[28, Definition 6.2] An $R$-$S$-homomorphism $f\!:A\to B$ from an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ to an $S$$\mathop{\widehat{\otimes}}$-algebra $B$ is called a two-sided relative homological epimorphism if $f$ is an epimorphism of $\mathop{\widehat{\otimes}}$-algebras and $A$ is acyclic relative to the functor
@@
We omit the definitions of left and right relative homological epimorphisms; for details see [28, Definition 6.1] with corrections in [29, 30]. The only fact that we need here is that a two-sided relative homological epimorphism is also left and right [28, Proposition 6.5].
@@
see [34, Definition 1.3], [27, Definition 3.2] or [28, Definition 6.3].
Lemma 4.3.
Every homological epimorphism from a finitely generated algebra to a Fréchet algebra satisfies Property (UDE).
Proof.
It follows from the definitions that every homological epimorphism is $1$-pseudoflat in the sense of [10]. The argument for the claim that a $1$-pseudoflat epimorphism from a finitely generated to a Fréchet algebra satisfies Property (UDE) is exactly the same as for [10, Theorem 3.24] (in the corrected version of [11]), where the case of a $1$-pseudoflat epimorphism between Fréchet algebras is considered; see also details in [12]. ∎
@@
For the proof we need a result on homological epimorphisms for analytic smash products with the Hopf $\mathop{\widehat{\otimes}}$-algebra $\mathfrak{A}_{s}$ defined in (2.7); see Theorem 4.6 below.
@@
The following proposition is the first step.
@@
Since $\mathfrak{A}_{s}$ is a Hopf $\mathop{\widehat{\otimes}}$-algebra by Lemma 2.4, it follows from [27, Proposition 3.7] that it suffices to check that $\mathbb{C}[z]\to\mathfrak{A}_{s}$ is a weak localization, i.e., the following conditions hold:
(1) for some projective resolution $0\leftarrow\mathbb{C}\leftarrow P_{\bullet}$ in $\mathbb{C}[z]\mbox{-}\!\mathop{\mathsf{mod}}$ the complex $0\leftarrow\mathfrak{A}_{s}\mathop{\widehat{\otimes}}_{\mathbb{C}[z]}\mathbb{C}\leftarrow\mathfrak{A}_{s}\mathop{\widehat{\otimes}}_{\mathbb{C}[z]}P_{\bullet}$ is admissible;
@@
Condition (2) is clearly satisfied. To check Condition (1) note that
@@
where $\varepsilon\!:z\mapsto 0$ and $[\delta_{0}(a)](z)\!:=az$, is a projective resolution in $\mathbb{C}[z]\mbox{-}\!\mathop{\mathsf{mod}}$. Applying $\mathfrak{A}_{s}\mathop{\widehat{\otimes}}_{\mathbb{C}[z]}(-)$ to this resolution we get
@@
Thus Condition (1) is also satisfied and therefore $\mathbb{C}[z]\to\mathfrak{A}_{s}$ is a weak localization for every $s$. ∎

### v4

@@ 2. Hopf $\mathop{\widehat{\otimes}}$-algebras, analytic smash products and Arens–Michael envelopes
Recall that an Arens–Michael algebra is a complete topological algebra whose topology can be determined by a family of submultiplicative seminorms. The Arens–Michael envelope of a topological algebra $A$ is a pair $(\widehat{A},\iota_{A})$, where $\widehat{A}$ is an Arens–Michael algebra and $\iota_{A}$ is a continuous homomorphism $A\to\widehat{A}$ such that for every Arens–Michael algebra $B$ and every continuous homomorphism $\varphi\!:A\to B$ there is a unique continuous homomorphism $\widehat{\varphi}\!:\widehat{A}\to B$ making the diagram
@@
commutative. More concretely, we can take as $\iota_{A}$ the completion homomorphism with respect to the topology determined by all possible continuous submultiplicative seminorms on $A$.
In the Lie algebra case, we can replace $U(\mathfrak{g})$ by $\mathfrak{g}$ in the diagram (assuming that $\varphi$ is a Lie algebra homomorphism) and say that $\widehat{U}(\mathfrak{g})$ is the Arens–Michael envelope of $\mathfrak{g}$.
@@
Let $\mathfrak{g}$ be a finite-dimensional solvable complex Lie algebra. Then there is an iterated semidirect sum decomposition
$\mathfrak{g}=((\cdots(\mathfrak{f}_{1}\rtimes\mathfrak{f}_{2})\rtimes\cdots)\rtimes\mathfrak{f}_{n},$ (2.8)
where $\mathfrak{f}_{1},\ldots,\mathfrak{f}_{n}$ are $1$-dimensional, and a non-increasing sequence $i_{1},\ldots,i_{n}$ in $[0,\infty]$ such that
@@
We use reduction to algebras of analytic functionals. Let $G$ be a simply connected complex Lie group whose Lie algebra is isomorphic to $\mathfrak{g}$. Also, let $\mathscr{A}(G)$ denote the corresponding algebra of analytic functionals (the strong dual space of the space holomorphic functions on $G$ endowed with the convolution multiplication) and $\widehat{\mathscr{A}}(G)$ the Arens–Michael envelope of $\mathscr{A}(G)$. Then the natural embedding $U(\mathfrak{g})\to\mathscr{A}(G)$ induces a topological isomorphism $\widehat{U}(\mathfrak{g})\to\widehat{\mathscr{A}}(G)$ [2, Proposition 2.1]. So we can apply results on $\widehat{\mathscr{A}}(G)$, Theorems 4.4, 6.5 and 6.4 in [9], which give the existence of the decomposition in (2.8) and (2.9), and the compatibility with the decomposition of $U(\mathfrak{g})$, respectively. ∎
@@
We use the results in this section in the proof of Theorem 3.5.
@@
$h\cdot\varphi(\alpha(r))=\sum\psi(h_{(1)})\,\varphi(\alpha(r))\,\psi(S(h_{(2)}))=\chi\left(\sum j(h_{(1)})\,i_{R}(r)\,j(S(h_{(2)}))\right)=\chi(i_{R}(h\cdot r)).$
(The last equality follows from the fact that $i_{R}$ is an $H$$\mathbin{\widehat{\otimes}}$-module morphism with respect to the adjoint action.) Thus $h\cdot\varphi(\alpha(r))=\varphi(\alpha(h\cdot r))$. Hence $\varphi\alpha$ is an $H$$\mathbin{\widehat{\otimes}}$-module morphism.
Further, by the hypothesis, $\alpha$ is also an $H$$\mathbin{\widehat{\otimes}}$-module morphism and so is $\varphi$ since the range of $\alpha$ is dense. Hence, by the universal property of smash products written for $S\mathop{\widehat{\#}}H$, there is a unique $\mathop{\widehat{\otimes}}$-algebra homomorphism $\theta$ such that $\theta i_{S}=\varphi$ and $\theta(\alpha\mathop{\widehat{\#}}1)j=\psi$.
Since $\varphi\alpha$ is an $H$$\mathbin{\widehat{\otimes}}$-module algebra homomorphism, it follows from the universal property of smash products written for $R\mathop{\widehat{\#}}H$ that a homomorphism $\chi$ such that $\chi i_{R}=\varphi\alpha$ and $\chi j=\psi$ is unique. Therefore $\theta(\alpha\mathop{\widehat{\#}}1)=\chi$.
@@
Hence, by the universal property of smash products written for $S\mathop{\widehat{\#}}K$, there is a unique $\mathop{\widehat{\otimes}}$-algebra homomorphism $\theta$ such that $\theta j_{K}=\psi$ and $\theta(1\mathop{\widehat{\#}}\beta)i=\varphi$. The first equality implies that $\theta(1\mathop{\widehat{\#}}\beta)j_{H}=\psi\beta$.
Being a $K$$\mathbin{\widehat{\otimes}}$-module morphism, $\varphi$ is also an $H$$\mathbin{\widehat{\otimes}}$-module morphism. Then the universal property of smash products written for $S\mathop{\widehat{\#}}H$ implies that a homomorphism $\chi$ such that $\chi i=\varphi$ and $\chi j_{H}=\psi\beta$ is unique. Therefore $\theta(1\mathop{\widehat{\#}}\beta)=\chi$.
@@ 3. Unique extension property for derivations
We denote the vector space of continuous derivations from a $\mathop{\widehat{\otimes}}$-algebra $A$ to an $A$$\mathop{\widehat{\otimes}}$-bimodule $X$ by $\operatorname{Der}(A,X)$. It is clear that a $\mathop{\widehat{\otimes}}$-algebra homomorphism $\varphi\!:A\to B$ induces the linear map
$\widetilde{\varphi}_{X}\!:\operatorname{Der}(B,X)\to\operatorname{Der}(A,X)\!:D\mapsto D\varphi.$
@@
It is proved in [26] that every Arens–Michael envelope satisfies Property (UDE).
In what follows we denote the pushout of $\mathop{\widehat{\otimes}}$-algebra homomorphisms $A\to B_{1}$ and $A\to B_{2}$ using the relative free algebra notation, i.e., as $B_{1}\mathbin{\ast_{A}}B_{2}$.
@@
The category of $\mathop{\widehat{\otimes}}$-algebras is cocomplete.
Proof.
It suffices to show that coproducts and coequalizers of pairs exist. The first assertion is proved in [8, Proposition B.1 and the remark after it] for finite coproducts. In the general case the argument is similar.
To prove the second assertion suppose that $\varphi,\psi\!:A\to B$ are parallel $\mathop{\widehat{\otimes}}$-algebra homomorphisms. Denote by $C$ the completion of the quotient $B/I$, where $I$ is a closed two-sided ideal generated by $\{\varphi(a)-\psi(a);\,a\in A\}$. (We need the completion since the quotient space may be non-complete.)
We claim that $(C,\tau)$, where $\tau\!:B\to C$ is the naturally defined homomorphism, is a coequalizer. Indeed, let $\chi\!:B\to D$ be a $\mathop{\widehat{\otimes}}$-algebra homomorphism such that $\chi\varphi=\chi\psi$. Then $\chi(\varphi(a)-\psi(a))=0$ for every $a$ and so $\chi$ maps $I$ to 00. Hence $\chi$ induces a continuous homomorphism $\chi^{\prime}\!:B/I\to D$ such that $\chi=\chi^{\prime}\tau$. Since $\chi^{\prime}$ extends continuously to $C$, the claim is proved. ∎
In particular, pushouts of $\mathop{\widehat{\otimes}}$-algebras exist. The following result is an analytic version of Proposition 5.2 in [12].
@@
In the category of unital $\mathop{\widehat{\otimes}}$-algebras, Property (UDE) is preserved by pushouts, i.e., if in a pushout diagram
[…]
the homomorphism $\varphi_{1}$ satisfies Property (UDE), then $\varkappa_{2}$ so does.
For the proof we need an auxiliary proposition. Note that every $(B_{1}\mathbin{\ast_{A}}B_{2})$$\mathop{\widehat{\otimes}}$-bimodule is both a $B_{1}$$\mathop{\widehat{\otimes}}$- and $B_{2}$$\mathop{\widehat{\otimes}}$-bimodule with the multiplications given by restriction of the canonical homomorphisms $\varkappa_{1}$ and $\varkappa_{2}$.
Proposition 3.4.
Let $\varphi_{1}\!:A\to B_{1}$ and $\varphi_{2}\!:A\to B_{2}$ be $\mathop{\widehat{\otimes}}$-algebra homomorphisms and $X$ be a $(B_{1}\mathbin{\ast_{A}}B_{2})$$\mathop{\widehat{\otimes}}$-bimodule. Suppose that $D_{1}\in\operatorname{Der}(B_{1},X)$, $D_{2}\in\operatorname{Der}(B_{2},X)$ and $D_{1}\varphi_{1}=D_{2}\varphi_{2}$. Then there exists a unique continuous derivation $D$ making the diagram
@@
$(c,x)(c^{\prime},x^{\prime})=(cc^{\prime},cx^{\prime}+xc^{\prime})\qquad(c,c^{\prime}\in B_{1}\mathbin{\ast_{A}}B_{2},\;x,x^{\prime}\in X).$
@@
Consider the pushout diagram in (3.1) and suppose that $\varphi_{1}$ satisfies Property (UDE). Take a $(B_{1}\ast_{A}B_{2})$$\mathop{\widehat{\otimes}}$-bimodule $X$ and $D_{2}\in\operatorname{Der}(B_{2},X)$. Since $D_{2}\varphi_{2}$ is a derivation of $A$ and $\varphi_{1}$ satisfies Property (UDE), there exists $D_{1}\in\operatorname{Der}(B_{1},X)$ such that $D_{1}\varphi_{1}=D_{2}\varphi_{2}$. By the universal property for derivations in Proposition 3.4, there is $D\in\operatorname{Der}(B_{1}\ast_{A}B_{2},X)$ such that $D_{1}=D\varkappa_{1}$ and $D_{2}=D\varkappa_{2}$. Hence, $\operatorname{Der}(B_{1}\ast_{A}B_{2},X)\to\operatorname{Der}(B_{2},X)$ is surjective.
Note that $\widetilde{\varphi}_{X}$ is injective for each $X$ if and only if $\varphi$ is an epimorphism; cf. the Fréchet algebra case in [10, Theorem 3.20]. Since epimorphisms are preserved by pushouts in every category, $\operatorname{Der}(B_{1}\ast_{A}B_{2},X)\to\operatorname{Der}(B_{2},X)$ is injective for each $X$. Thus, $\varkappa_{2}$ satisfies Property (UDE). ∎
@@
It is easy to see that Property (UDE) is stable under composition. So it suffices to show that $\alpha\mathop{\widehat{\#}}1$ and $1\mathop{\widehat{\#}}\beta$ have this property. Since Property (UDE) is preserved by pushouts according to Proposition 3.3, the result follows from Propositions 2.6 and 2.7, which assert that $\alpha\mathop{\widehat{\#}}1$ and $1\mathop{\widehat{\#}}\beta$ are obtained by pushouts with $\alpha$ and $\beta$, respectively. ∎
@@ 4. Homological epimorphisms and relatively quasi-free algebras
Suppose that $R$ is a $\mathop{\widehat{\otimes}}$-algebra. Recall that an $R$$\mathop{\widehat{\otimes}}$-algebra is a pair $(A,\eta_{A})$, where $A$ is a $\mathop{\widehat{\otimes}}$-algebra and $\eta_{A}\!:R\to A$ is a $\mathop{\widehat{\otimes}}$-algebra homomorphism. Note that each $A$$\mathop{\widehat{\otimes}}$-module is automatically an $R$$\mathop{\widehat{\otimes}}$-module via the restriction-of-scalars functor along $\eta_{A}$. We denote by $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}$ the category of left $A$$\mathop{\widehat{\otimes}}$-modules endowed with the split exact structure relative to $R$. This means that the admissible sequences in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}$ are those that are split by $R$$\mathop{\widehat{\otimes}}$-module morphisms; cf. [26, Appendix, Example 10.1 and 10.3]. In particular, when $R=\mathbb{C}$, we recover the standard definition of an admissible (or $\mathbb{C}$-split) sequence of $A$$\mathop{\widehat{\otimes}}$-modules used in [17]. When considering $\mathop{\widehat{\otimes}}$-bimodules over an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ (from the left) and an $S$$\mathop{\widehat{\otimes}}$-algebra $B$ (from the right), we denote the corresponding category by $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(B,S)$. In the case when $R=S=\mathbb{C}$, we write simply $A\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}B$. When we talk on a projective resolution in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(B,S)$, we mean a complex consisting of objects projective in the corresponding exact category and splitting in $R\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}S$.
Let $A$ be a $\mathop{\widehat{\otimes}}$-algebra and let $X$ and $Y$ be a right and left $A$$\mathop{\widehat{\otimes}}$-module, respectively. Then the projective $A$-module tensor product $X\mathop{\widehat{\otimes}}_{A}Y$ is defined as the completion of the quotient space of $X\mathop{\widehat{\otimes}}Y$ by the closure of the linear hull of all elements of the form $x\cdot a\otimes y-x\otimes a\cdot y$ ($x\in X$, $y\in Y$, $a\in A$); see [17].
@@
[26, Definition 6.2] An $R$-$S$-homomorphism $f\!:A\to B$ from an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ to an $S$$\mathop{\widehat{\otimes}}$-algebra $B$ is called a two-sided relative homological epimorphism if $f$ is an epimorphism of $\mathop{\widehat{\otimes}}$-algebras and $A$ is acyclic relative to the functor
@@
We omit the definitions of left and right relative homological epimorphisms; for details see [26, Definition 6.1] with corrections in [28, 29]. The only fact that we need here is that a two-sided relative homological epimorphism is also left and right [26, Proposition 6.5].
@@
see [31, Definition 1.3], [25, Definition 3.2] or [26, Definition 6.3].
@@
For the proof we need a result on homological epimorphisms for analytic smash products with the Hopf $\mathop{\widehat{\otimes}}$-algebra $\mathfrak{A}_{s}$ defined in (2.7) but in a more general formulation; see Theorem 4.5 below.
@@
The following proposition is the first step of the proof of Theorem 4.3.
@@
Since $\mathfrak{A}_{s}$ is a Hopf $\mathop{\widehat{\otimes}}$-algebra by Lemma 2.4, it follows from [25, Proposition 3.7] that it suffices to check that $\mathbb{C}[z]\to\mathfrak{A}_{s}$ is a weak localization, i.e., the following conditions hold:
(1) for some projective resolution $0\leftarrow\mathbb{C}\leftarrow P_{\bullet}$ in $\mathbb{C}[z]\mbox{-}\!\mathop{\mathsf{mod}}$ the complex
$0\leftarrow\mathfrak{A}_{s}\mathop{\widehat{\otimes}}_{\mathbb{C}[z]}\mathbb{C}\leftarrow\mathfrak{A}_{s}\mathop{\widehat{\otimes}}_{\mathbb{C}[z]}P_{\bullet}$
is admissible;
@@
To check Condition (1) note that
@@
where $\varepsilon\!:z\mapsto 0$ and $[d_{0}(a)](z)\!:=az$, is a projective resolution in $\mathbb{C}[z]\mbox{-}\!\mathop{\mathsf{mod}}$. Applying $\mathfrak{A}_{s}\mathop{\widehat{\otimes}}_{\mathbb{C}[z]}(-)$ to this resolution we get
@@
To check Condition (2) recall that for a $\mathop{\widehat{\otimes}}$-algebra $A$, a right Fréchet module $X$, and a closed ideal $I$ such that $A/I$ is also a Fréchet space the well-known formula $X\mathop{\widehat{\otimes}}_{A}(A/I)\cong X/\,\overline{X\cdot I}$ holds; see, e.g., [27, Proposition 3.1]. When $A=\mathbb{C}[z]$, $X=\mathfrak{A}_{s}$ and $I$ is the ideal of polynomials vanishing at 00, we have the desired isomorphism $\mathfrak{A}_{s}\mathop{\widehat{\otimes}}_{\mathbb{C}[z]}\mathbb{C}\to\mathbb{C}$.
Thus both condition are satisfied and therefore $\mathbb{C}[z]\to\mathfrak{A}_{s}$ is a weak localization for every $s$. ∎